If 3.41 g of salt is separated from 5.69 g of a mixture of salt and said, what is the percentage of salt and sand in the mixture. show all work
Answer #1

Since 5.69 g of mixture has 3.41 g of salt

Then 1 g of mixture will have amount of salt  

= 3.41/5.69  

Hence 100 g of mixture will have amount of salt ( percent of salt)

= 100×3.41/5.69

= 59.9 %

Now amount of sand = 5.69 - 3.41  

= 2.28 g

Percent of sand in mixture = 2.28×100/5.69

= 40.1 %

Or,

Percent of sand = 100 - percent of salt

= 100 - 59.9

= 40.1 %